A body of mass 10 kg is moving with a constant velocity of 10 m/s. When a constant force acts for 4 seconds on it, it moves with a velocity 2 m/sec in the opposite direction. The acceleration produced in it is. A. 3m/s^2. B. -3m/s^2. C.0.3m/s^2. D. -0.3m/s^2

Grade:

2 Answers

Saurabh Singh
10 years ago
VELOCITY CHANGE = (10-(-2))
= 12 M/S

TIME PERIOD = 4 SEC

ACCELERATION = 12/4
= 3 M/S^2

HENCE A

Thanks & Regards

Saurabh Singh,

askIITians Faculty

B.Tech.

IIT Kanpur

Sunil Kumar FP
10 years ago
inital velocity u=10m/s2
final velocity v=-2m/s2
we know that v=u+ at
-2=10+a×4
a=-3m/s2

thanks and regards
sunil kr
askIItian faculty